The complete combustion of octane with oxygen, represented by the unbalanced
chemical equation: C8H18 + O2 → CO2 + H2O
10 grams of octane and 10grams of oxygen will yield what mass of water?
1
Expert's answer
2020-06-22T09:14:43-0400

Solution:

The balanced chemical equation:

2C8H18 + 25O2 → 16CO2 + 18H2

According to the chemical equation: n(C8H18)/2 = n(O2)/25 = n(H2O)/18


Convert all given information into moles:

Moles of octane = n(C8H18) = Mass of C8H18 / Molar mass of C8H18

The molar mass of C8H18 is 114 g/mol

n(C8H18) = (10 g) / (114 g/mol) = 0.0877 mol

Moles of oxygen = n(O2) = Mass of O2 / Molar mass of O2

The molar mass of O2 is 32 g/mol

n(O2) = (10 g) / (32 g/mol) = 0.3125 mol


Divide by coefficients of balanced equation:

octane ⇒ 0.0877 mol / 2 mol = 0.04385

oxygen ⇒ 0.3125 mol / 25 mol = 0.0125

Oxygen is the lower value. It is the limiting reactant.


We now use the limiting reactant to find the amount of water produced.

n(O2)/25 = n(H2O)/18

n(H2O) = 18 × n(O2) / 25

n(H2O) = 18 × (0.3125 mol) / 25 = 0.225 mol


Moles of H2O = n(H2O) = Mass of H2O / Molar mass of H2O

The molar mass of H2O is 18 g/mol.

Finally, the mass of H2O is:

Mass of H2O = (0.225 mol) × (18 g/mol) = 4.05 g


Answer: 4.05 grams of water will be produce.